kaylak:

calc help Get an equation for the radius of the red balloon as a function of time. (Hint: Drag the slider to t=1, and then t=2, and then t=3, etc, noting each time what the radius of the Red Balloon is. See if you can spot the pattern, and come up with an equation. (the equation will look like: � = ___ �)

kaylak:

r=_t @Vocaloid

Vocaloid:

well just try plugging in several values of t and see what sort of output values you get

kaylak:

the radius increases by .2 each time so .2t

kaylak:

Write an equation for the Volume of a sphere as a function of time. Remember: The Volume of a Sphere is given by: � = ! ! ��! (i.e. substitute your answer from the previous question in for “�”)

kaylak:

so r is .2

kaylak:

I guess v(t)=4/3pi.2^3

Vocaloid:

r = 0.2t as determined by the previous step V = (4/3)pi * r^3 so you can just plug r = 0.2t into the volume equation

Vocaloid:

since it asks for V in terms of t, your equation should have t in it

kaylak:

this may be easier to read

1 attachment
kaylak:

okay

kaylak:

so the rate of change question wants the derivative of my equation

Vocaloid:

yup

kaylak:

isn't the derivative of a sphere like 4pi r ^2 or something

Vocaloid:

it wants it in terms of t not r from step b) you should have V = (4/3)pi * (0.2t)^3 take the derivative of this, with respect to t, to get c)

kaylak:

so dt/dv?

Vocaloid:

dV/dt

kaylak:

oh crud the rate changes after t=3

Vocaloid:

huh. weird.

kaylak:

nevermind

kaylak:

it doesn't thank god

kaylak:

2.513t^2

Vocaloid:

uh I think I got something a bit diff. V = (4/3)pi * (0.2t)^3 dV/dt = (4/3)(pi)(3)(0.2^3)t^2

Vocaloid:

I'd leave it in terms of pi

kaylak:

2 attachments
Vocaloid:

it's already in terms of V and t so just take the derivative

Vocaloid:

then for e) set the two V'(t) equations equal to each other and solve for t

kaylak:

is it dv/dt again?

kaylak:

I got 10

Vocaloid:

yes, good

kaylak:

e blue 10 20 30 40 etc red .03 .27 .9 2.14 4.19 7.24 11.49 17.16 24.43 33.51

Vocaloid:

you just need to solve dV/dt red = dV/dt blue dV/dt red = (4/3)(pi)(3)(0.2^3)t^2 = 0.032pi * t^2 then just set that equal to 10 and solve for t.

kaylak:

so plug in 10 at the equal sign?

Vocaloid:

dV/dt red = 0.032pi * t^2 dV/dt blue = 10 therefore 0.032pi * t^2 = 10

kaylak:

is that it ?

kaylak:

g is last question

Vocaloid:

like I said you need to solve for t.

kaylak:

oh crap okay

Vocaloid:

for g, you simply create two volume equations green and orange will both be V = (4/3)pi*r^3 then plug in the given expression for r into each equation, to generate two separate equations for V in terms of t, one for each balloon

kaylak:

t is about 10 it's not 9.8 let me see if it is 9.9

Vocaloid:

then set the derivatives equal to each other

Vocaloid:

I end up w/ t≈9.97356 for e

kaylak:

what did you do to get your answer 9.9 etc?

Vocaloid:

0.032pi * t^2 = 10 t^2 = 10/0.032pi t = sqrt(10/0.032pi)

kaylak:

green (4/3)pi*10+5t^3?

Vocaloid:

careful with your parentheses (4/3)pi*[10+5t]^3?

kaylak:

orange same idea with[]?

kaylak:

so what volume would the 2 increase at the same rate

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid then set the derivatives equal to each other \(\color{#0cbb34}{\text{End of Quote}}\)

Vocaloid:

set the derivatives equal to each other and solve for t

kaylak:

so are the equations I made the derivative or do I have t find it? slightly confused

Vocaloid:

the equations the volume equations in terms of t. you will then take the derivatives.

kaylak:

20pi/3

kaylak:

for green

kaylak:

gotta go in a bit but I hope I am doing it right

Vocaloid:

check your work again V = (4/3)pi*(10+5t)^3 apply chain rule dV/dt = (4/3)(pi)*3*(10+5t)^2 * (5)

kaylak:

200piy(10+5t)

kaylak:

I am lost and being rushed so I might have to do this later

Vocaloid:

sure

kaylak:

I have a few minutes if you could help me at least get the green side

kaylak:

@Vocaloid

Vocaloid:

that is the green one if you're talking about the orange one V(t) = (4/3)pi*(2+t^2)^3 chain rule V'(t) = (4/3)(pi)(3)(2+t^2)^2 * 2t

Vocaloid:

at that point you just have to set the two derivatives equal to get (4/3)(pi)(3)(2+t^2)^2 * 2t = (4/3)(pi)*3*(10+5t)^2 * (5), solve for t

kaylak:

1 attachment
kaylak:

?

kaylak:

@Vocaloid

kaylak:

@Hero

kaylak:

@Vocaloid

kaylak:

the pic above is my answer is that correct

Vocaloid:

No, remember you are taking the derivative of both volume expressions

kaylak:

20pix(10+5t)^2 for green

Vocaloid:

Yes

kaylak:

really ?

Vocaloid:

Yes Now set it equal to the orange balloon derivative

kaylak:

I have 8pitx(2+t^2)^2 for orange

Vocaloid:

Good

kaylak:

is that it?

Vocaloid:

Again you have to set them equal to each other and solve for t

kaylak:

how do I do that like combine them like basic algebra?

Vocaloid:

Uh Solving it manually is kind of a pain because you have to expand everything and factor If possible I would recommend using a calculator

kaylak:

I literally got an error

kaylak:

I need this assignment done today could you just set the euation up for I can calculate it in terms of t

kaylak:

@Vocaloid

kaylak:

?

Vocaloid:

20pi*(10+5t)^2 = (4/3)(pi)(3)(2+t^2)^2 * 2t t≈4.7408